OpenStudy (anonymous):

If the length of each side of a cube is reduced to one-third its original value, is the surface area reduced to one-third its original value? Use an example to support your answer.

OpenStudy (anonymous):

finally i got the answer :P

OpenStudy (wolf1728):

Volume of a cube = (side)^3 Surface area of a cube = (side)^2 * 6 So, if we reduce one side to one third, then the surface area becomes (1/3)^2 * 6 (1/9) * 6 Original area = (1) * 6 New area = (1/9) * 6 So if we reduce one side to a third then the surface area is reduced to 1/9 of that it was originally. (Although you didn't ask this, when a side is reduced to 1/3 then the volume decreases by 1/27)